OpenStudy (anonymous):

SIMPLIFY \[\sqrt{108}-3\sqrt{8}-\sqrt{27}\]

OpenStudy (anonymous):

sqrt(3*4*9) - 3*sqrt(2*2*2) - sqrt (3*3*3); 2*3*sqrt(3) - 3*2*sqrt(2) - 3*sqrt (3); 6*sqrt (3) - 6*sqrt(2) - 3*sqrt(3); 3*sqrt(3) - 6*sqrt(2)

OpenStudy (maheshmeghwal9):

\[6\sqrt{3}-6\sqrt{2}-3\sqrt{3}.\]

OpenStudy (maheshmeghwal9):

\[\implies 3\sqrt{3}-6\sqrt{2}.\]